imqwerty (imqwerty):

in an example it was given - v^2=-10s+2500 2vdt/dt=-10ds/dt please explain the second step.

OpenStudy (irishboy123):

there's a typo in this (dt/dt) but we can do it anyway you start with \[v^2=-10s+2500 \] you are now going to differentiate with respect to t this requires *implicit* differentiation \[\frac{d \ (v^2)}{dt} = 2 \ v \ \frac{dv}{dt}\] \[\frac{d \ (-10s)}{dt} = -10 \ \frac{ds}{dt}\]

imqwerty (imqwerty):

thanks @IrishBoy123 :)

imqwerty:

Picklein typo in that book got me here